The sum of the squares of three consecutive even integers is 116. Find the integers.
Answer by checkley71(8403)

X~2+(X+2)~2+(X+4)~2=116 OR X~2+X~2+4X+4+X~2+8X+16=116 OR 3X~2+12X+20=116 OR
3X~2+12X-96=0 OR X~2+4X-32=0 OR (X+8)(X-4)=0 OR X+8=0 OR X=-8 & X-4=0 OR X=4
PROOF 4*4+6*6+8*8=116 OR 16+36+64=116 OR 116=116